# Nonlinear Permuted Granger Causality

### v1

@@ Abstract
Granger causal inference is a contentious but widespread method used in fields ranging from economics to neuroscience. The original definition addresses the notion of causality in time series by establishing functional dependence conditional on a specified model. Adaptation of Granger causality to nonlinear data remains challenging, and many methods apply in-sample tests that do not incorporate out-of-sample predictability leading to concerns of model overfitting. To allow for out-of-sample comparison, we explicitly define a measure of functional connectivity using permutations of the covariate set. Artificial neural networks serve as featurizers of the data to approximate any arbitrary, nonlinear relationship, and under certain conditions on the featurization process and the model residuals, we prove consistent estimation of the variance for each permutation. Performance of the permutation method is compared to penalized objective, naive replacement, and omission techniques via simulation, and we investigate its application to neuronal responses of acoustic stimuli in the auditory cortex of anesthetized rats. We contend that targeted use of the Granger causal framework, when prior knowledge of the causal mechanisms in a dataset are limited, can help to reveal potential predictive relationships between sets of variables that warrant further study.
@@ 1 Introduction
Granger causal inference investigates the ability of a set of variables $\mathbf{X}\in\mathbb{R}^{T\times p}$ to predict future values of a response $\mathbf{Y}\in\mathbb{R}^{T\times d}$ in a multivariate time series, $t=1,\ldots,T$ (Granger, 1969; Wiener, 1956). The effect is traditionally measured through the variance of residuals in restricted and unrestricted models, like shown in Definition 1, where $\mathcal{P}(\mathbf{y}_{t}|\mathcal{I}_{<t})$ represents the optimal prediction of vector $\mathbf{y}_{t}$ given all information prior to time $t$. Subscripts indicating time constraints of the set of variables $\mathbf{X}$ are not included as we trivially assume that the covariates assigned to a vector response $\mathbf{y}_{t}$ precede time $t$.
Definition 1.
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ is Granger causal for multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ if $\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\right)\right]<\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\backslash\mathbf{X}\right)\right]$.
1.1 Discussion of Granger Causality
@@
The form of Granger causality presented in Definition 1 is inherently conditional on additional information included in the modeling process. The optimal prediction $\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\right)$ is unattainable in practice, and the notion of Granger causality is a conditional model on some included explanatory covariate set $\mathbf{Z}\in\mathbb{R}^{T\times q}$ and the history of the response $\mathbf{Y}_{<t}$ like that in Definition 2.
Definition 2.
@@
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ is conditionally Granger causal for multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ given $\mathbf{Z}\in\mathbb{R}^{T\times q}$ and the relevant history of the response $\mathbf{Y}_{<t}$ if $\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X}\right)\right]<\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}\right)\right]$.
@@
Model selection is implicit in the test for presence of Granger causality, and the inferential conclusion is coupled with a written form (Friston et al., 2013). Inclusion of additional variables strengthens the condition for establishing Granger causality; rejection of the null implies the covariate set $\mathbf{X}$ is found to provide unique and useful information for prediction of $\mathbf{y}_{t}$ beyond that contained in both $\mathbf{Z}$ and the lagged response (Granger, 1980). The basis for Granger causality requires fulfillment of several conditions including a sufficient length of continuous-valued stationary data, exact and complete specification of the model, error-free observation of the variables, and a sampling frequency on a regular discrete grid that contains the known lag relationship (Granger, 1969, 1980, 1988). In the form of Definition 2, we can relax the second condition provided that inference is accordingly narrowed to the conditional statement.
@@
The framework can also be defined in terms of non-causality as a statement of conditional independence, where inclusion of additional variables $\mathbf{Z}$ is a simple extension (Florens and Mouchart, 1982; Granger, 1980).
Definition 3.
@@
Multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times p}$ does not Granger cause multivariate time series $\mathbf{Y}\in\mathbb{R}^{T\times d}$ if and only if $\mathbf{Y}_{<t+1}\perp\mathbf{X}$ given $\mathbf{Y}_{<t}$.
@@
This statement is mathematically equivalent, but perhaps more powerful because it is defined in terms of the distributions of the variables, allowing for extension to several other forms of statistical tests. This definition may be prone to misuse if interpreted to place the burden of proof on establishing independence. In a definition from Section 4 of Shojaie and Fox (2022), time series $\mathbf{x}_{j}$ is Granger non-causal for time series $\mathbf{x}_{i}$ if and only if $\forall t$, $\mathcal{P}\left(\mathbf{x}_{ti}|\mathbf{x}_{<t1},\ldots,\mathbf{x}_{<tj},\ldots,\mathbf{x}_{<tp}\right)=\mathcal{P}\left(\mathbf{x}_{ti}|\mathbf{x}_{<t1},\ldots,\mathbf{x}_{<t(j-1)},\mathbf{x}_{<t(j+1)},\ldots,\mathbf{x}_{<tp}\right)$, or the equality of these predictions holds for all time points $t$. If misinterpreted to mean finding correlation at one time point in series $\mathbf{x}_{i}$ is enough to claim functional dependence of two time series, the presence of a causal connection effectively becomes the null hypothesis and Granger causality is reduced to an exceptionally weak statement.
@@
Even when the question is framed with the onus on demonstrating a functional relationship, all Granger causal methods overreach their scope of reliable application when inference is performed on the individual variables included in the covariate set rather than on their collective behavior. Applications of the framework to interpreting individual model coefficients introduce a hidden multiplicity problem of repeated testing on subsets of $\mathbf{X}$, and the conclusion requires amendment to conditional non-causality of $\mathbf{x}_{i}$ given an exhaustive list of all other components $\mathbf{x}_{j}$, $j=1,,\ldots,i-1,i+1,\ldots,p$ in the model after adequate Type 1 error control. Methods that select causal covariate pairs through the use of penalized optimization problems don’t always allow for easy extension to the multiple testing problem, and may require repetitive sub-sampling approaches such as stability selection (Meinshausen and Bühlmann, 2010). A thorough definition of Granger causality provides a clear representation of the collective conclusion to be drawn on the covariate set $\mathbf{X}$, clarifies the conditional nature of the result on the specified model and included variables, and stresses the philosophical ordering from the null hypothesis implying no causal structure to the alternative that demands evidence of the contrary, all while retaining any general functional form of $\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X}\right)$.
@@
We note the predictive nature of these definitions can relate to a causal structure between two variable groups, but alone is not enough to establish effective connectivity, distinguishing a direct influence of one population on another (Bressler and Seth, 2011; Friston, 1994). Even in the presence of the optimal set $\mathcal{I}_{<t}$, association and precedence are not enough to distinguish true causality if slight redundancies are included or an effect does not remain constant in direction through time (Maziarz, 2015). Granger causality exists in the realm of functional connectivity that identifies correlation at one or more time lags (Friston, 1994). Appropriate use of Granger causality is contentious, but the method has been applied to a variety of fields like economics, environmental sciences, and neuroscience (Bernanke, 1990; Cox Jr and Popken, 2015; Holland, 1986; Reid et al., 2019; Seth et al., 2015; Sims, 1972). Cautious and targeted use of the Granger causal framework can elucidate predictive relationships between variables that warrant further study when prior knowledge of potential causal relationships is limited.
@@
In the linear realm, $\mathcal{P}(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z},\mathbf{X})$ is often sought from a vector autoregression (VAR) model, where models are evaluated with in-sample tests (Granger, 1969, 1980). Geweke (1982) proposed a spectral decomposition form of linear Granger causality for application to stationary Gaussian processes. Inference is performed with the estimated covariance of the restricted model (with $\mathbf{X}$ excluded) and that of the unrestricted model, where under the null it is assumed the two are equal. Increasing dimension of the response variable often requires implementation of an approximate test or a switch to permutation-like testing (Anderson and Robinson, 2001; Barnett and Seth, 2011). In-sample testing differs from the true notion of predictive ability, and out-of-sample methods align closer to the essence of Granger causality (Chao et al., 2001; Inoue and Kilian, 2005; Peters et al., 2016). This distinction is especially important when using deep learning techniques to model complex, nonlinear dynamics.
@@ 1.2 Nonlinear Adaptations
Nonparametric methods provide the basis for many nonlinear adaptations of Granger causality; specification of the exact functional form can be challenging. Parametric attempts, like the ordinary differential equations approaches of Henderson and Michailidis (2014) and Wu et al. (2014), allow for flexible definition of a series of functions to capture dependence, but are limited to modelling additive dynamics when the true mechanism of interaction may be more complicated. Some model-free information theoretic methods detect more elaborate forms of nonlinear dependence with minimal assumptions, but suffer from highly variable estimates and challenges of application to multivariate systems (Amblard and Michel, 2011; Runge et al., 2012; Vicente et al., 2011). Kernel Granger causality examines the linear form in a transformed feature space, but model comparison can become difficult (Marinazzo et al., 2008, 2011). Artificial neural networks allow for general forms of nonlinear dependence in a similar feature space.
@@
Fully connected feed-forward networks (with $g$ a sigmoid function), like the construction shown in Equations 1 through 3, fall under the umbrella of the universal approximation theorems of Cybenko (1989) and Hornik (1991). These artificial networks act as featurizers of the data to a higher dimensional space, $\Psi:\mathbf{X}\in\mathbb{R}^{T\times p}\rightarrow\mathbf{H}\in\mathbb{R}^{T\times N}$, $N\gg p$; the architecture allows for arbitrarily close approximation $\hat{\mathbf{y}}_{t}$ of any function output $\mathbf{y}_{t}=f(\mathbf{x}_{t})$, provided $\mathbf{h}_{t}\in\mathbb{R}^{N}$ is allowed to contain a sufficient width of hidden states $N$ (Barron, 1993).
$\mathbf{h}^{1}_{t}$ $=g\left(\mathbf{W}^{1}\mathbf{x}_{t}+\mathbf{b}^{1}\right)$ (1)
$\mathbf{h}^{\ell}_{t}$ $=g\left(\mathbf{W}^{\ell}\mathbf{h}^{\ell}_{t}+\mathbf{b}^{\ell}\right)$ (2)
$\hat{\mathbf{y}}_{t}$ $=\mathbf{W}^{L}\mathbf{h}_{t}+\mathbf{b}^{L}$ (3)
@@
In Equations 1 and 2, $g$ is an element-wise activation function, and $\ell=1,\ldots,L$ is the number of layers, or the “depth” of the network, where $\mathbf{W}^{1}\in\mathbb{R}^{N\times p}$, $\mathbf{W}^{\ell}\in\mathbb{R}^{N\times N}$, $\mathbf{W}^{L}\in\mathbb{R}^{d\times N}$, $\mathbf{b}^{\ell}\in\mathbb{R}^{N}$, and $\mathbf{b}^{L}\in\mathbb{R}^{d}$ are parameter matrices. The artificial networks shown above, sometimes called multilayer perceptrons (MLPs), and recurrent neural networks (RNNs) like long-short term memory networks (LSTMs) (Graves, 2012) and echo state networks (ESNs) (Jaeger, 2007) have all been used to investigate functional connectivity in time series data (Shojaie and Fox, 2022). RNN-type network construction adds a recurrence term to the hidden state form in Equation 1, $\mathbf{h}_{t}=g\left(\mathbf{W}^{r}\mathbf{h}_{t-1}+\mathbf{W}^{1}\mathbf{x}_{t}+\mathbf{b}\right)$.
@@
Tank et al. (2022) extend Granger causality to the nonlinear space using component-wise MLP (cMLP) models, which model individual variables in the response $\mathbf{y}_{i}$, $i=1,\ldots,d$, with separate artificial networks. Parameters are sought via a penalized optimization approach and proximal gradient descent, and no Granger causal connection is inferred for an individual covariate if the corresponding row in a component input parameter matrix $\mathbf{W}^{1}_{i}$ is zero (Tank et al., 2022). The penalized approach encourages sparse solutions that block the inclusion of information from less predictive components in the hidden states $\mathbf{h}_{t}$, but selecting the regularization parameter is not an easy task and values may produce vastly different results. Tank et al. (2022) further introduce a component-wise LSTM (cLSTM) model that harnesses the recurrent structure to circumvent selection of the optimal lag for inclusion in the covariate set $\mathbf{X}$. This formulation, while making model specification as it relates to the time lag components easier, has the consequence of muddling inferential results across several lags. Khanna and Tan (2019) build on the framework of cMLP with statistical recurrent units, Biswas and Ombao (2022) discuss the application of the component network structure to frequency-specific relationships and non-stationary data, and Marcinkevics and Vogt (2021) introduce generalized vector autoregressive (GVAR) methodology aimed at interpretability of potential functional relationships. Other methods like the Jacobian Granger causality (JGC) method of Suryadi et al. (2023) use the Jacobian matrix, and Nauta et al. (2019) use convolutional neural networks (TCDF) and attention scores to serve as measures of variable importance.
@@
Jointly estimating a large number of parameters is computationally expensive, and Duggento et al. (2021) instead use randomly initialized ESNs. The matrices $\mathbf{W}^{1}$ and $\mathbf{b}^{1}$ are populated with random values, and learning only takes place in the output weights $\mathbf{W}^{L}$ via penalized regression (Jaeger, 2007). Because computation is performed using linear techniques rather than a gradient descent algorithm, complexity decreases; however, inference can only be performed on the output coefficients if information mixing does not occur in the hidden states $\mathbf{h}_{t}$. They formulate $\mathbf{W}^{1}$ as a block diagonal matrix, which limits the scope of application to a specific subset of additive nonlinear interactions (Duggento et al., 2021).
@@
Many of these methods ignore the collective inference principle of Granger causality and instead take the eager approach of performing individual covariate inference, sometimes with disregard for the multiplicity problem. Evaluation of these predictive relationships is often performed using in-sample tests. With their universal approximator ability, artificial neural networks of sufficient width or depth can approximate any functional relationship between two covariate sets, even if it is data-specific and the model is overfit, making them prone to link variables that do not have a predictive relationship as the dimension of the network increases. Sparsity inducing penalties may marginally improve reliability of in-sample tests, but out-of-sample testing helps control the overfitting problem to identify only useful functional relationships.
In this vain, Horvath et al. (2022) develop the Learned Kernel VAR (LeKVAR) method that proposes use of a kernel parameterized by an artificial neural network they argue is less prone to overfitting from a decoupling importance measure of the individual series and the selected lags, and the TCDF method employs a permutation-like testing procedure after variable selection (Nauta et al., 2019). Our goal is to formulate a simple method for identification of nonlinear functional relationships that lends itself to out-of-sample testing and clear control of the multiplicity problem in a Granger causal framework.
1.3 Overview

### v2

@@ Abstract
Granger causal inference is a contentious but widespread method used in fields ranging from economics to neuroscience. The original definition addresses the notion of causality in time series by establishing functional dependence conditional on a specified model. Adaptation of Granger causality to nonlinear data remains challenging, and many methods apply in-sample tests that do not incorporate out-of-sample predictability, leading to concerns of model overfitting. To allow for out-of-sample comparison, a measure of functional connectivity is explicitly defined using permutations of the covariate set. Artificial neural networks serve as featurizers of the data to approximate any arbitrary, nonlinear relationship, and consistent estimation of the variance for each permutation is shown under certain conditions on the featurization process and the model residuals. Performance of the permutation method is compared to penalized variable selection, naive replacement, and omission techniques via simulation, and it is applied to neuronal responses of acoustic stimuli in the auditory cortex of anesthetized rats. Targeted use of the Granger causal framework, when prior knowledge of the causal mechanisms in a dataset are limited, can help to reveal potential predictive relationships between sets of variables that warrant further study.
@@ 1 Introduction
Granger causal inference investigates the ability of a time series $\mathbf{x}_{t}\in\mathbb{R}^{p}$, $t=1,\ldots,T$, to predict future values of a response $\mathbf{y}_{t}\in\mathbb{R}^{d}$ (Wiener, 1956; Granger, 1969). The effect is traditionally measured through the variance of residuals in restricted and unrestricted models, like shown in Definition 1.1, where $\mathcal{P}$ represents the optimal prediction function, $\mathcal{I}_{<t}$ is all information prior to time $t$, and $\mathbf{X}_{<t}$ is a matrix of compiled values of $\mathbf{x}_{t}$ prior to time $t$.
Definition 1.1.
Time series $\mathbf{x}_{t}\in\mathbb{R}^{p}$ is Granger causal for $\mathbf{y}_{t}\in\mathbb{R}^{d}$ if
$\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\right)\right]<\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\backslash\mathbf{X}_{<t}\right)\right].$ (1)
Modern methods for adapting Granger causality to nonlinear functional relationships leverage deep learning and representation learning for capturing dependence between variables. These tools, when brazenly paired with other machine learning techniques, are not necessarily reliable or precise. Penalized variable selection as a screening method forces the isolation of information; this does not consider the collective covariate set of the system and tends to up-weight contributions of the chosen nonzero covariates. Erroneous conclusions can result from deep learning when variable specific inference is muddled. This work explicitly redefines Granger causality in terms of a permuted framework with out-of-sample testing (NPGC) that retains the flexibility of representation learning and has specific advantages when seeking nonlinear functional connections.
1.1 Granger Causality
@@
The form of Granger causality presented in Definition 1.1 is inherently conditional on additional information included in the modeling process. The optimal prediction $\mathcal{P}\left(\mathbf{y}_{t}|\mathcal{I}_{<t}\right)$ is unattainable in practice, and the notion of Granger causality is a conditional model on some included explanatory covariate set $\mathbf{z}_{t}\in\mathbb{R}^{q}$ and the history of the response prior to time $t$, $\mathbf{Y}_{<t}$, like that in Definition 1.2.
Definition 1.2.
@@
Time series $\mathbf{x}_{t}\in\mathbb{R}^{p}$ is conditionally Granger causal for $\mathbf{y}_{t}\in\mathbb{R}^{d}$ given $\mathbf{z}_{t}\in\mathbb{R}^{q}$ and the relevant history of the response if
$\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}_{<t},\mathbf{X}_{<t}\right)\right]<\text{Var}\left[\mathbf{y}_{t}-\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}_{<t}\right)\right].$ (2)
@@
Model selection is implicit in the test for presence of Granger causality, and the inferential conclusion is coupled with a written form (Friston et al., 2013). Inclusion of additional variables strengthens the condition for establishing Granger causality; rejection of the null implies the covariate set $\mathbf{x}_{t}$ is found to provide unique and useful information for prediction of $\mathbf{y}_{t}$ beyond that contained in both $\mathbf{z}_{t}$ and the lagged response (Granger, 1980). The basis for Granger causality requires fulfillment of several conditions including a sufficient length of continuous-valued stationary data, exact and complete specification of the model, error-free observation of the variables, and a sampling frequency on a regular discrete grid that contains the known lag relationship (Granger, 1969, 1980, 1988). In the form of Definition 1.2, the second condition can be relaxed provided that inference is accordingly narrowed to the conditional statement.
@@
The framework can also be defined in terms of non-causality as a statement of conditional independence, where inclusion of additional variables $\mathbf{z}_{t}$ is a simple extension (Granger, 1980; Florens and Mouchart, 1982).
Definition 1.3.
@@
Time series $\mathbf{x}_{t}\in\mathbb{R}^{p}$ does not Granger cause $\mathbf{y}_{t}\in\mathbb{R}^{d}$ if and only if
$\mathbf{Y}_{<t+1}\perp\mathbf{X}_{<t}\quad\text{given}\quad\mathbf{Y}_{<t}.$ (3)
@@
This statement is equivalent, but perhaps more powerful because it is defined in terms of the distributions of the variables, allowing for extension to several other forms of statistical tests. It may be prone to misuse if interpreted to place the burden of proof on establishing independence. In a definition from Section 4 of Shojaie and Fox (2022), column $j$ in time series $\mathbf{x}_{t}$ is Granger non-causal for time series $\mathbf{y}_{t}$ if and only if $\forall t$,
$\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{x}_{<t1},\ldots,\mathbf{x}_{<tj},\ldots,\mathbf{x}_{<tp}\right)=\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{x}_{<t1},\ldots,\mathbf{x}_{<t(j-1)},\mathbf{x}_{<t(j+1)},\ldots,\mathbf{x}_{<tp}\right)$ (4)
that implies the equality of these predictions holds for all time points $t$. If misinterpreted to mean finding correlation at one time point in series $\mathbf{y}_{t}$ is enough to claim functional dependence of two time series, the presence of a causal connection effectively becomes the null hypothesis and Granger causality is reduced to an exceptionally weak statement.
@@
Even when the question is framed with the onus on demonstrating a functional relationship, all Granger causal methods overreach their scope of reliable application when inference is performed on the individual variables included in the covariate set rather than on their collective behavior. Applications of the framework to interpreting individual model coefficients introduce a hidden multiplicity problem of repeated testing on subsets of $\mathbf{x}_{t}$, and the conclusion requires amendment to conditional non-causality of $\mathbf{y}_{t}$ given an exhaustive list of all other components $\mathbf{x}_{tj}$, $j=1,\ldots,p$ in the model after adequate Type 1 error control. Methods that select causal covariate pairs through the use of penalized optimization problems don’t always allow for easy extension to the multiple testing problem, and may require repetitive sub-sampling approaches such as stability selection (Meinshausen and Bühlmann, 2010). A thorough definition of Granger causality provides a clear representation of the collective conclusion to be drawn on the covariate set $\mathbf{x}_{t}$, clarifies the conditional nature of the result on the specified model and included variables, and stresses the philosophical ordering from the null hypothesis implying no causal structure to the alternative that demands evidence of the contrary, all while retaining any general functional form of $\mathcal{P}\left(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}_{<t},\mathbf{X}_{<t}\right)$.
@@
Holland (1986) relates the definition to that of Suppes (1970) and criticizes its fragile reliance on the specified pre-exposure variables that may completely change an inferential result. Maziarz (2015) writes that “Granger causality does not meet the requirements of an investigator who uses this method due to epistemic reasons” and the methodology should be used “only if the theoretical background is insufficient,” noting the common cause fallacy, indirect causality, and problems related to sampling frequency. In this tone, the predictive nature of these definitions can relate to a causal structure between two variable groups, but alone is not enough to establish effective connectivity, distinguishing a direct influence of one population on another (Bressler and Seth, 2011; Friston, 1994). Even in the presence of the optimal set $\mathcal{I}_{<t}$, association and precedence are not enough to distinguish true causality if slight redundancies are included or an effect does not remain constant in direction through time (Maziarz, 2015). Granger causality exists in the realm of functional connectivity that identifies correlation at one or more time lags (Friston, 1994). Appropriate use of Granger causality is contentious, but the method has been applied to a variety of fields like economics, environmental sciences, and neuroscience (Bernanke, 1990; Cox Jr. and Popken, 2015; Holland, 1986; Reid et al., 2019; Seth et al., 2015; Sims, 1972). Cautious and targeted use of the Granger causal framework can elucidate predictive relationships between variables that warrant further study when prior knowledge of potential causal relationships is limited.
@@
In the linear realm, $\mathcal{P}(\mathbf{y}_{t}|\mathbf{Y}_{<t},\mathbf{Z}_{<t},\mathbf{X}_{<t})$ is often sought from a VAR model and evaluated with in-sample testing (Granger, 1969, 1980). Geweke (1982) proposed a spectral decomposition form of linear Granger causality for application to stationary Gaussian processes. Inference is performed with the estimated covariance of the restricted model (with $\mathbf{X}_{<t}$ excluded) and that of the unrestricted model, where under the null it is assumed the two are equal. Increasing dimension of the response variable often requires implementation of an approximate test or a switch to permutation-like testing (Anderson and Robinson, 2001; Barnett and Seth, 2011). In-sample testing differs from the true notion of predictive ability, and out-of-sample methods align closer to the essence of Granger causality (Chao et al., 2001; Inoue and Kilian, 2005; Peters et al., 2016). This distinction is especially important when using deep learning techniques to model complex, nonlinear dynamics.
@@ 1.2 Nonlinear Adaptations
Nonparametric methods provide the basis for many nonlinear adaptations of Granger causality; specification of the exact functional form can be challenging. Parametric attempts, like the ordinary differential equations approaches of Henderson and Michailidis (2014) and Wu et al. (2014), allow for flexible definition of a series of functions to capture dependence, but are limited to modeling additive dynamics when the true mechanism of interaction may be more complicated. Some model-free information theoretic methods detect more elaborate forms of nonlinear dependence with minimal assumptions, but suffer from highly variable estimates and challenges of application to multivariate systems (Amblard and Michel, 2011; Runge et al., 2012; Vicente et al., 2011). Kernel Granger causality examines the linear form in a transformed feature space, but model comparison can become difficult (Marinazzo et al., 2008, 2011). Artificial neural networks (ANNs) allow for general forms of nonlinear dependence in a similar feature space.
@@
Fully connected feed-forward networks (with $g$ a sigmoid function), like the construction shown in Equations 5 through 7, fall under the umbrella of the universal approximation theorems of Cybenko (1989) and Hornik (1991). These artificial networks act as featurizers of the data to a high-dimensional space, $\Psi:\mathbf{X}\in\mathbb{R}^{T\times p}\rightarrow\mathbf{H}\in\mathbb{R}^{T\times N}$, $N\gg p$; the architecture allows for arbitrarily close approximation $\hat{\mathbf{y}}_{t}$ of any function output $\mathbf{y}_{t}=f(\mathbf{x}_{t})$, provided $\mathbf{h}_{t}\in\mathbb{R}^{N}$ is allowed to contain a sufficient width of hidden states $N$.
$\mathbf{h}^{1}_{t}$ $=g\left(\mathbf{W}^{0}\mathbf{x}_{t}+\mathbf{b}^{0}\right)$ (5)
$\mathbf{h}^{\ell+1}_{t}$ $=g\left(\mathbf{W}^{\ell}\mathbf{h}^{\ell}_{t}+\mathbf{b}^{\ell}\right)$ (6)
$\hat{\mathbf{y}}_{t}$ $=\mathbf{W}^{L}\mathbf{h}^{L}_{t}+\mathbf{b}^{L}$ (7)
@@
In Equations 5 and 6, $g$ is an element-wise activation function, and $\ell=1,\ldots,L$ is the number of layers, or the “depth” of the network, where $\mathbf{W}^{0}\in\mathbb{R}^{N\times p}$, $\mathbf{W}^{\ell}\in\mathbb{R}^{N\times N}$, $\mathbf{W}^{L}\in\mathbb{R}^{d\times N}$, $\mathbf{b}^{0},\mathbf{b}^{\ell}\in\mathbb{R}^{N}$, and $\mathbf{b}^{L}\in\mathbb{R}^{d}$ are parameter matrices. The artificial networks shown above, sometimes called multilayer perceptrons (MLPs), and recurrent neural networks (RNNs) like long-short term memory networks (LSTMs) (Hochreiter and Schmidhuber, 1997; Graves, 2012) and echo state networks (ESNs) (Jaeger, 2007) have all been used to investigate functional connectivity in time series data (Shojaie and Fox, 2022). RNN-type network construction adds a recurrence term to the hidden state form in Equations 5 and 6, $\mathbf{h}^{1}_{t}=g\left(\mathbf{W}^{1h}\mathbf{h}^{1}_{t-1}+\mathbf{W}^{0}\mathbf{x}_{t}+\mathbf{b}^{0}\right)$, and $\mathbf{h}^{\ell+1}_{t}=g\left(\mathbf{W}^{\ell h}\mathbf{h}^{\ell+1}_{t-1}+\mathbf{W}^{\ell}\mathbf{h}^{\ell}_{t}+\mathbf{b}^{\ell}\right)$.
@@
Tank et al. (2022) extend Granger causality to the nonlinear space using component-wise MLPs (cMLP), which model individual variables in the response $\mathbf{y}_{ti}$, $i=1,\ldots,d$, with separate artificial networks. Parameters are sought via a penalized optimization approach and proximal gradient descent, and no Granger causal connection is inferred for an individual covariate if the corresponding row in a component input parameter matrix ($\mathbf{W}^{0}_{i}$ in Equation 5) is zero (Tank et al., 2022). The penalized approach encourages sparse solutions that block the inclusion of information from less predictive components in the hidden states $\mathbf{h}_{t}$, but selecting the regularization parameter is not an easy task and values may produce vastly different results. Tank et al. (2022) further introduce a component-wise LSTM (cLSTM) model that harnesses the recurrent structure to circumvent selection of the optimal lag for inclusion in the covariate set $\mathbf{X}_{<t}$. This formulation, while making model specification as it relates to the time lag components easier, has the consequence of mixing inferential results across several lags.
@@
Other methods for capturing arbitrary, nonlinear functional relationships include Khanna and Tan (2019) that builds on the framework of cMLP with statistical recurrent units, Biswas and Ombao (2022) that discusses the application of the component network structure to frequency-specific relationships and non-stationary data, and Marcinkevičs and Vogt (2021) that introduces generalized vector autoregressive (GVAR) methodology aimed at interpretability of potential functional relationships. The Jacobian Granger causality method of Suryadi et al. (2023) uses the Jacobian matrix, and Nauta et al. (2019) (TCDF) uses convolutional neural networks (CNNs) and attention scores to serve as measures of variable importance. Jointly estimating a large number of parameters is computationally expensive, and Duggento et al. (2021) instead use randomly initialized ESNs. Because computation is performed using linear techniques rather than a gradient descent algorithm, complexity decreases; however, inference can only be performed on the output coefficients if information mixing does not occur in the hidden states $\mathbf{h}_{t}$. They formulate each $\mathbf{W}^{h}$ in the RNN structure as a block diagonal matrix, which limits the scope of application to a specific subset of additive nonlinear interactions (Duggento et al., 2021).
@@
Many of these methods ignore the collective inference principle of Granger causality and instead take the eager approach of performing individual covariate inference, sometimes with disregard for the multiplicity problem. Evaluation of these predictive relationships is often performed using in-sample tests. With their universal approximator ability, artificial neural networks of sufficient width or depth can approximate any functional relationship between two covariate sets, even if it is data-specific and the model is overfit, making them prone to link variables that do not have a predictive relationship as the dimension of the network increases. Sparsity inducing penalties may marginally improve reliability of in-sample tests, but out-of-sample testing helps control the overfitting problem to identify only useful functional relationships. In this vein, Horvath et al. (2022) develop the Learned Kernel VAR (LeKVAR) method that proposes use of a kernel parameterized by an artificial neural network they argue is less prone to overfitting from a decoupling importance measure of the individual series and the selected lags, and the TCDF method employs a permutation-like testing procedure after variable selection (Nauta et al., 2019).